which of the accerlation vs time graphs is correct for a book sliding across a horizontal frictionsless table in 1-d at a constant velocity of 3.0 m/s?

ty of 3.0 m/s? acceleration, respectively
Answer #1

Solution) The area under acceleration time graph gives change in velocity

Here it is mentioned that velocity is constant that is 3m/s

So acceleration should be zero as acceleration a = dv/dt

a = d(3)/dt = 0

So from given graphs it should be graph D